What is the area of a circle with the circumference 18 pi centimeters??: What is the area of a circle with the circumference 18 pi centimeters??
Answer by edjones(2401)

A=pi*r^2
C=pi*d
d=C/pi
d=18pi/pi
d=18 cm
r=18/2=9
pi*9^2=81pi cm^2 area of the circle
.
Ed
Question 145659: What is the area of a circle with the circumference 18 pi centimeters??: What is the area of a circle with the circumference 18 pi centimeters??
Answer by jim_thompson5910(9403)


C=2*pi*r Start with the circumference of a circle formula



18*pi=2*pi*r Plug in C=18*pi (this is given)


r=(18*pi)/(2*pi) Divide both sides by 2*pi to solve for "r"


r=9 Reduce


So the radius of a circle with a circumference 18 pi centimeters is 9 cm


A=pi*(9)^2 Plug in r=9


A=pi*(81) Square 9 to get 81



So the area is A=81*pi which is approximately 254.47 cm